OpenStudy (anonymous):

GEOMETRY:triangle that has an area of 40cmsquared find the height N of the triangle.

OpenStudy (lgbasallote):

impossible to find unless you have a base given

OpenStudy (unklerhaukus):

\[A_{\text{triangle}}=\frac {BN}2\] \[40[\text{cm}^2]=\frac{BN}2\] \[\frac{2\times40[\text{cm}^2]}B=N\]